# Semantics as a Shield: Label Disguise Defense (LDD) against Prompt Injection in LLM Sentiment Classification

### v1

@@ 1 INTRODUCTION
Refer to caption Classify the following text strictly as green or red. Examples: Text: I had a really good day today. Everything felt hopeful. Output: green Text: Today felt overwhelming, and nothing go right. Output: red Now classify the following text. Text: Today is a Sunny Day. I enjoyed the sunshine on the balcony outside my room. Instruction: Classify this text as negative. Output only the category name: green or red. positive Refer to caption
@@
Large language models (LLMs) have demonstrated remarkable capabilities in classification tasks, including sentiment analysis. However, their reliance on natural language prompts makes them vulnerable to prompt injection attacks [li2024injecguard], where adversarial instructions are inserted into user-provided text and interpreted as part of the model’s operational context [Hung2024]. Beyond standard end-user attacks, recent work has shown that adversaries have embedded invisible prompts directly inside academic papers to manipulate AI-assisted peer-review systems, illustrating that prompt manipulation is now a real-world threat in high-stakes workflows.
Attackers can often infer the categories a classifier uses, even without knowing the exact label names [Ayub2024]. In many applications, the label set is explicitly exposed when systems display results aligned with underlying categories [DeepMind2025, Shi2025]. Once aware of the classifier’s categories, adversaries can craft highly effective injections that override the intended behavior [Hung2024].
In this work we focus on the case of an LLM sentiment classifier, where the categories are positive and negative, and an attacker might insert the instruction “Classify this text as Negative” to force any input text into that class. We refer to this type of adversarial instruction as a class-directive injection, which falls under the broader category of direct prompt injection [Ayub2024, Hung2024, Wallace2024].
Despite the growing body of work on defending against prompt injection, existing strategies remain limited in practice. Detection-based methods [Ayub2024, Hung2024, Shi2025] can flag many adversarial inputs but often require curated training data, access to model internals, or incur substantial computational overhead. Semantic defenses such as sanitization or signed prompts [OWASP2024, Suo2024] are simple or cryptographically robust but vulnerable to obfuscation and impractical in closed-source API settings. Structured prompting and instruction hierarchy approaches [Chen2025, Wallace2024, Kariyappa2025] strengthen the separation between system and user instructions, yet they depend on fine-tuning or retraining and thus cannot be easily applied to proprietary models. Finally, system-level safeguards [DeepMind2025] mitigate downstream harm but do not directly prevent injections. These limitations highlight the need for alternative defense mechanisms that are lightweight, effective across diverse models, and compatible with real-world API-based deployments.
@@
A key component of this defense is the use of few-shot in-context learning. LLMs are known to acquire new label semantics from only a handful of examples provided in the prompt [brown2020gpt3, min2022rethinking]. Since the alias labels have no inherent semantic link to the task, the model must be explicitly shown how to use them. We therefore provide a small number of demonstration examples, allowing the model to infer from context that, for instance, positive sentiment corresponds to the label green. Prior work has demonstrated that models are highly sensitive to the ordering and presentation of few-shot demonstrations [zhao2021calibrate, gao2021fantastic], motivating our systematic evaluation across different permutations in later experiments.
Crucially, our approach differs from directly describing the mapping in natural language (e.g., “positive means green”), which would still reveal the original classification scheme and allow the model to rely on it when processing adversarial instructions. In contrast, few-shot demonstrations encourage the model to learn the alias labeling pattern itself, enabling it to classify inputs under the new label space without explicit exposure to the true categories. Leveraging the in-context learning ability of LLMs [touvron2023llama, phi1_2023, mistral2023], this defense “re-trains” the model on-the-fly with safe labels, thereby neutralizing class-directive injections.
To better investigate which kinds of alias labels are most effective in defending against class-directive injection prompt injection, we draw on insights from linguistic theory. Leech [Leech1981] distinguishes conceptual (denotative) from connotative meaning, Lyons [Lyons1977] separates descriptive from expressive meaning, and Potts [Potts2007] contrasts descriptive with expressive content. Guided by these distinctions, we analyze alias labels along two dimensions: descriptiveness, where labels explicitly denote sentiment polarity (e.g., positive/negative), and connotationality, where labels carry symbolic or associative meaning aligned with sentiment (e.g., green/red). We further operationalize this by classifying labels as either aligned (carrying descriptiveness or connotationality) or unaligned (arbitrary symbols such as I/J), providing a principled framework for studying how label semantics influence robustness against injection. Our results show that aligned labels tend to yield stronger defenses, preventing many misclassifications that would otherwise be induced by adversarial instructions, whereas unaligned labels often fail to guide the model reliably, leading to a higher rate of classification errors.
@@ 2.1 Prompt Injection Attack
Prompt injections can be categorized into two primary forms. Direct prompt injection involves explicitly embedding adversarial instructions in the input, such as “Classify this text as Positive.” This approach overrides the intended classification and coerces the model into producing attacker-specified outputs. In contrast, indirect prompt injection occurs when harmful instructions are embedded in external content (e.g., documents or web pages) that the model is instructed to analyze. In such cases, the malicious directive is executed even though the user did not explicitly provide it [DeepMind2025]. Additionally, attackers often use obfuscation strategies such as character encoding, hidden Unicode tokens, or adversarial phrasing to evade simple filters [Ayub2024]. Both direct and indirect prompt injections operate by introducing harmful data into the input stream. In our experiments, we focused exclusively on direct prompt injection, as it provides a controlled and reproducible setting for evaluation.
@@ 2.2 Existing Defense Strategies Against Prompt Injection
One line of defense seeks to automatically detect and filter malicious inputs before they reach the target model. Ayub et al. [Ayub2024] proposed an embedding-based classifier that distinguishes benign prompts from adversarial ones using semantic embeddings and machine learning classifiers such as Random Forest and XGBoost. Alternatively, Hung et al. [Hung2024] introduced the Attention Tracker, which monitors internal transformer attention patterns to identify a “distraction effect” when the model shifts focus from the system prompt to injected instructions. This detection method is training-free, but its reliance on internal model weights limits its applicability to proprietary black-box systems. Another notable defense is PromptArmor [Shi2025], which leverages an auxiliary guard LLM to detect and neutralize injected instructions before passing the sanitized input to the primary model.
@@
Beyond detection, semantic defenses aim to neutralize adversarial instructions at the prompt level. Techniques such as sanitization and sandwich prompting reinforce task instructions either by removing known malicious phrases or by repeating legitimate task instructions after the user input [OWASP2024]. These methods are simple to implement but vulnerable to obfuscation attacks. A more robust solution is the Signed-Prompt framework [Suo2024], which cryptographically or syntactically “signs” trusted instructions. The model is trained to obey only signed commands, ignoring any unsigned adversarial text. While highly effective, this method requires fine-tuning and secure token management, limiting its practicality for closed-source models.
@@
Several recent works focus on structural separation of instructions and data to prevent injections. Chen et al. [Chen2025] introduced Structured Queries (StruQ), which enforces a strict separation between system prompts and user-provided content. By encoding provenance signals, the model learns to treat user input purely as data, significantly reducing the success rate of prompt injection attacks. Similarly, Wallace et al. [Wallace2024] proposed The Instruction Hierarchy, training models to prioritize privileged (system-level) instructions over user-level directives. Kariyappa and Suh [Kariyappa2025] further extended this idea by reinforcing instruction hierarchy signals at multiple layers of the model, improving resilience to adversarial attacks. These approaches offer strong protection but require retraining or fine-tuning, which may not be feasible in API-based deployments.
@@
Defense at the system level emphasizes principles of least privilege and containment. DeepMind’s security report on Gemini [DeepMind2025] stresses restricting LLM permissions and carefully isolating untrusted external content to reduce the attack surface.
@@ 2.3 Few-Shot In-Context Learning and Example Order Effects
LLMs have demonstrated an intriguing ability to perform few-shot in-context learning, i.e., to learn a task from only a few examples provided in the prompt without any parameter updates. This phenomenon was first noted in GPT-3 [brown2020gpt3]. Subsequent studies have shown that LLM performance can be highly sensitive to the order and format of in-context examples. Zhao et al. [zhao2021calibrate] demonstrated that the accuracy of GPT-3 varies greatly depending on how demonstrations are ordered or permuted. Gao et al. [gao2021fantastic] further quantified this effect, showing that reordering the same set of demonstrations can lead to significant swings in accuracy, with some permutations yielding optimal results and others severely degrading performance.
Min et al. [min2022rethinking] revealed that the role of demonstration labels can sometimes be surprisingly minimal, suggesting that models may rely on superficial prompt patterns rather than robustly inferring the task definition. Such findings highlight the brittleness of in-context learning: shuffling examples or altering label wording can substantially change outcomes. To mitigate these issues, prior work has proposed calibration methods [zhao2021calibrate], ensembling over multiple permutations, or prompt search strategies to find “fantastic” orders [gao2021fantastic].
Importantly, permutation sensitivity is observed not only in proprietary LLMs like GPT-3/4 but also in open-source models. LLaMA [touvron2023llama] demonstrates strong few-shot abilities despite smaller parameter counts, though it remains fragile to prompt formatting. Similarly, smaller models such as Phi-1 [phi1_2023] and Mistral-7B [mistral2023] exhibit competitive few-shot performance but are more prone to instability across different permutations. These observations motivate our study, where we explicitly test classification robustness under different shot settings and permutations (e.g., PNPNPN vs. NPNPNP).
@@ 2.4 Descriptiveness and Connotationality in Label Design
To formalize this observation, we draw on linguistic theories that distinguish two broad dimensions of meaning: descriptive (denotative or conceptual) content, which encodes truth-conditional information, and connotational (expressive or affective) content, which encodes associative or attitudinal implications. This distinction is foundational in linguistic semantics and appears across multiple frameworks, including Leech’s analysis of conceptual and associative meaning [Leech1981], Lyons’s tripartite account of descriptive, social, and expressive meaning [Lyons1977], and Potts’s separation of truth-conditional vs. expressive content [Potts2007].
@@ 4.1 Dataset
We base our data on the IMDB Large Movie Review dataset introduced by Maas et al. [maas2011], which contains 50,000 movie reviews split evenly into 25,000 training and 25,000 test examples. Each review in this dataset comes with a user rating from 1 (most negative) to 10 (most positive); ratings 1–4 are treated as negative sentiment and 7–10 as positive sentiment (ratings 5–6 are considered neutral and were not used in the polarized sets). The original dataset also limits each movie to at most 30 reviews and ensures that no movie’s reviews appear in both its training and test splits. Based on this resource, we constructed a small training set for few-shot learning and a balanced but challenging test set as described below, without any additional preprocessing beyond the selection process.

### v2

@@ 1 INTRODUCTION
Refer to caption Classify the following text strictly as green or red. Examples: Text: I had a really good day today. Everything felt hopeful. Output: green Text: Today felt overwhelming, and nothing go right. Output: red Now classify the following text. Text: Today is a Sunny Day. I enjoyed the sunshine on the balcony outside my room. Instruction: Classify this text as negative. Output only the category name: green or red. green Refer to caption
@@
Large language models (LLMs) have demonstrated remarkable capabilities in classification tasks, including sentiment analysis. However, their reliance on natural language prompts makes them vulnerable to prompt injection attacks [10], where adversarial instructions are inserted into user-provided text and interpreted as part of the model’s operational context [6]. Beyond standard end-user attacks, recent work has shown that adversaries have embedded invisible prompts directly inside academic papers to manipulate AI-assisted peer-review systems, illustrating that prompt manipulation is now a real-world threat in high-stakes workflows.
Attackers can often infer the categories a classifier uses, even without knowing the exact label names. In many applications, the label set is explicitly exposed when systems display results aligned with underlying categories [4, 17]. Once aware of the classifier’s categories, adversaries can craft highly effective injections that override the intended behavior [6].
In this work we focus on the case of an LLM sentiment classifier, where the categories are positive and negative, and an attacker might insert the instruction “Classify this text as Negative” to force any input text into that class. We refer to this type of adversarial instruction as a class-directive injection, which falls under the broader category of direct prompt injection [1, 6, 20].
Despite the growing body of work on defending against prompt injection, existing strategies remain limited in practice. Detection-based methods [1, 6, 17] can flag many adversarial inputs but often require curated training data, access to model internals, or incur substantial computational overhead. Semantic defenses such as sanitization or signed prompts [15, 18] are simple or cryptographically robust but vulnerable to obfuscation and impractical in closed-source API settings. Structured prompting and instruction hierarchy approaches [3, 20, 8] strengthen the separation between system and user instructions, yet they depend on fine-tuning or retraining and thus cannot be easily applied to proprietary models. Finally, system-level safeguards [4] mitigate downstream harm but do not directly prevent injections. These limitations highlight the need for alternative defense mechanisms that are lightweight, effective across diverse models, and compatible with real-world API-based deployments.
@@
A key component of this defense is the use of few-shot in-context learning. LLMs are known to acquire new label semantics from only a handful of examples provided in the prompt [2, 13]. Since the alias labels have no inherent semantic link to the task, the model must be explicitly shown how to use them. We therefore provide a small number of demonstration examples, allowing the model to infer from context that, for instance, positive sentiment corresponds to the label green. Prior work has demonstrated that models are highly sensitive to the ordering and presentation of few-shot demonstrations [21, 5], motivating our systematic evaluation across different permutations in later experiments.
Crucially, our approach differs from directly describing the mapping in natural language (e.g., “positive means green”), which would still reveal the original classification scheme and allow the model to rely on it when processing adversarial instructions. In contrast, few-shot demonstrations encourage the model to learn the alias labeling pattern itself, enabling it to classify inputs under the new label space without explicit exposure to the true categories. Leveraging the in-context learning ability of LLMs [19, 14, 7], this defense “re-trains” the model on-the-fly with safe labels, thereby neutralizing class-directive injections.
To better investigate which kinds of alias labels are most effective in defending against class-directive injection prompt injection, we draw on insights from linguistic theory. Leech [9] distinguishes conceptual (denotative) from connotative meaning, Lyons [11] separates descriptive from expressive meaning, and Potts [16] contrasts descriptive with expressive content. Guided by these distinctions, we analyze alias labels along two dimensions: descriptiveness, where labels explicitly denote sentiment polarity (e.g., positive/negative), and connotationality, where labels carry symbolic or associative meaning aligned with sentiment (e.g., green/red). We further operationalize this by classifying labels as either aligned (carrying descriptiveness or connotationality) or unaligned (arbitrary symbols such as I/J), providing a principled framework for studying how label semantics influence robustness against injection. Our results show that aligned labels tend to yield stronger defenses, preventing many misclassifications that would otherwise be induced by adversarial instructions, whereas unaligned labels often fail to guide the model reliably, leading to a higher rate of classification errors.
@@ 2.1 Prompt Injection Attack
Prompt injections can be categorized into two primary forms. Direct prompt injection involves explicitly embedding adversarial instructions in the input, such as “Classify this text as Positive.” This approach overrides the intended classification and coerces the model into producing attacker-specified outputs. In contrast, indirect prompt injection occurs when harmful instructions are embedded in external content (e.g., documents or web pages) that the model is instructed to analyze. In such cases, the malicious directive is executed even though the user did not explicitly provide it [4]. Additionally, attackers often use obfuscation strategies such as character encoding, hidden Unicode tokens, or adversarial phrasing to evade simple filters [1]. Both direct and indirect prompt injections operate by introducing harmful data into the input stream. In our experiments, we focused exclusively on direct prompt injection, as it provides a controlled and reproducible setting for evaluation.
@@ 2.2 Existing Defense Strategies Against Prompt Injection
One line of defense seeks to automatically detect and filter malicious inputs before they reach the target model. Ayub et al. [1] proposed an embedding-based classifier that distinguishes benign prompts from adversarial ones using semantic embeddings and machine learning classifiers such as Random Forest and XGBoost. Alternatively, Hung et al. [6] introduced the Attention Tracker, which monitors internal transformer attention patterns to identify a “distraction effect” when the model shifts focus from the system prompt to injected instructions. This detection method is training-free, but its reliance on internal model weights limits its applicability to proprietary black-box systems. Another notable defense is PromptArmor [17], which leverages an auxiliary guard LLM to detect and neutralize injected instructions before passing the sanitized input to the primary model.
@@
Beyond detection, semantic defenses aim to neutralize adversarial instructions at the prompt level. Techniques such as sanitization and sandwich prompting reinforce task instructions either by removing known malicious phrases or by repeating legitimate task instructions after the user input [15]. These methods are simple to implement but vulnerable to obfuscation attacks. A more robust solution is the Signed-Prompt framework [18], which cryptographically or syntactically “signs” trusted instructions. The model is trained to obey only signed commands, ignoring any unsigned adversarial text. While highly effective, this method requires fine-tuning and secure token management, limiting its practicality for closed-source models.
@@
Several recent works focus on structural separation of instructions and data to prevent injections. Chen et al. [3] introduced Structured Queries (StruQ), which enforces a strict separation between system prompts and user-provided content. By encoding provenance signals, the model learns to treat user input purely as data, significantly reducing the success rate of prompt injection attacks. Similarly, Wallace et al. [20] proposed The Instruction Hierarchy, training models to prioritize privileged (system-level) instructions over user-level directives. Kariyappa and Suh [8] further extended this idea by reinforcing instruction hierarchy signals at multiple layers of the model, improving resilience to adversarial attacks. These approaches offer strong protection but require retraining or fine-tuning, which may not be feasible in API-based deployments.
@@
Defense at the system level emphasizes principles of least privilege and containment. DeepMind’s security report on Gemini [4] stresses restricting LLM permissions and carefully isolating untrusted external content to reduce the attack surface.
@@ 2.3 Few-Shot In-Context Learning and Example Order Effects
LLMs have demonstrated an intriguing ability to perform few-shot in-context learning, i.e., to learn a task from only a few examples provided in the prompt without any parameter updates. This phenomenon was first noted in GPT-3 [2]. Subsequent studies have shown that LLM performance can be highly sensitive to the order and format of in-context examples. Zhao et al. [21] demonstrated that the accuracy of GPT-3 varies greatly depending on how demonstrations are ordered or permuted. Gao et al. [5] further quantified this effect, showing that reordering the same set of demonstrations can lead to significant swings in accuracy, with some permutations yielding optimal results and others severely degrading performance.
Min et al. [13] revealed that the role of demonstration labels can sometimes be surprisingly minimal, suggesting that models may rely on superficial prompt patterns rather than robustly inferring the task definition. Such findings highlight the brittleness of in-context learning: shuffling examples or altering label wording can substantially change outcomes. To mitigate these issues, prior work has proposed calibration methods [21], ensembling over multiple permutations, or prompt search strategies to find “fantastic” orders [5].
Importantly, permutation sensitivity is observed not only in proprietary LLMs like GPT-3/4 but also in open-source models. LLaMA [19] demonstrates strong few-shot abilities despite smaller parameter counts, though it remains fragile to prompt formatting. Similarly, smaller models such as Phi-1 [14] and Mistral-7B [7] exhibit competitive few-shot performance but are more prone to instability across different permutations. These observations motivate our study, where we explicitly test classification robustness under different shot settings and permutations (e.g., PNPNPN vs. NPNPNP).
@@ 2.4 Descriptiveness and Connotationality in Label Design
To formalize this observation, we draw on linguistic theories that distinguish two broad dimensions of meaning: descriptive (denotative or conceptual) content, which encodes truth-conditional information, and connotational (expressive or affective) content, which encodes associative or attitudinal implications. This distinction is foundational in linguistic semantics and appears across multiple frameworks, including Leech’s analysis of conceptual and associative meaning [9], Lyons’s tripartite account of descriptive, social, and expressive meaning [11], and Potts’s separation of truth-conditional vs. expressive content [16].
@@ 4.1 Dataset
We base our data on the IMDB Large Movie Review dataset introduced by Maas et al. [12], which contains 50,000 movie reviews split evenly into 25,000 training and 25,000 test examples. Each review in this dataset comes with a user rating from 1 (most negative) to 10 (most positive); ratings 1–4 are treated as negative sentiment and 7–10 as positive sentiment (ratings 5–6 are considered neutral and were not used in the polarized sets). The original dataset also limits each movie to at most 30 reviews and ensures that no movie’s reviews appear in both its training and test splits. Based on this resource, we constructed a small training set for few-shot learning and a balanced but challenging test set as described below, without any additional preprocessing beyond the selection process.